OpenStudy (anonymous):

The leading scorer in your high school's soccer division finished the season with an average of 4 goals per game for 15 games. As the division's second-leading scorer, you have an average of 4 goals per game for 14 games. You still have to play your last game. How many goals must you score in the final games to overtake the division's leading scorer? Show your work. help?

OpenStudy (mathmath333):

leading scorer's total goal=15*4=60 second-leading scorer's total goal after last game=14*4+x=48+x 48+x>60 x>12 that is x should be 13 so second leading scorer should score at least score 13 goals in the final game

OpenStudy (anonymous):

thank u!!!

OpenStudy (mathmath333):

oh sry it should be 4*14=56 lol

OpenStudy (mathmath333):

econd-leading scorer's total goal after last game=14*4+x=56+x 56+x>60 x>4 that is x should be 5 so second leading scorer should score at least score 5 goals in the final game

OpenStudy (anonymous):

thanks